what is the pH of 0.02 mole HCl added to one liter of acetic acid and sodium acetate


1
Expert's answer
2021-05-12T06:30:30-0400

Mole of acetic acid = molarity×volume= 1mol

Mole of sodium acetate= 1×1=1

If we add 0.02 mil HCl then salt will react .

After addition of salt

Mole of salt= 1-0.02=0.08

Mole of acid= 1.02

pH=pKa+log [salt]/[acid]

pH= 4.75+log[0.08/1.02]

pH=3.67